OpenStudy (anonymous):

HOW INTEGERATE THIS: y=6^x

OpenStudy (lgbasallote):

you integrate it...by deriving it..does that make sense/

hartnn (hartnn):

seriously? integral of a^x=a^x/ln a

OpenStudy (anonymous):

just remember how to differentiate it\[y=a^x\]\[\ln y=x \ln a\]take derivative\[\frac{y'}{y}=\ln a\]\[y'=y\ln a\]\[y'=a^x \ln a\]

OpenStudy (anonymous):

@hartnn u sure........ I got derivation of y=a^x ----> a^x * lna

OpenStudy (anonymous):

But I am asking for integration

hartnn (hartnn):

thats derivative,u asked integral a^x/ln a + c

OpenStudy (anonymous):

If I differentiate y=a^x/lna will I get a^x

hartnn (hartnn):

sure.y=a^x/lna+c u will get it.

hartnn (hartnn):

but diff. w.r.t x

OpenStudy (anonymous):

Can u PLZ SHOW it

hartnn (hartnn):